OpenStudy (anonymous):

I Will Medal And Fan Best Answer! I need explanation AND answer! Divide. Write your answer in simplest form. 5/8 divided 3/10

Vocaloid (vocaloid):

when we divide by fractions, we take the reciprocal of the second fraction and multiply (5/8)/(3/10) = (5/8)*(10/3) = ?

OpenStudy (anonymous):

hm. i have no clue what to do with ( )